# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 4.3 Exact reconstruction with MMA
    Assume $S\not\subseteq L[I]$. In that case, there must be a facet of codirection $j$, for some $j\in\mathcal{J}^{-}$, that intersects $S$. Since one has $c^{-}_{j}\leq(c^{-}+\overrightarrow{v})_{j}<\alpha_{j}$ for any $j\in\mathcal{J}^{-}$, this means that the facet would not intersect $K$, which yields to a contradiction as per Definition 4.4, item (4).
This concludes that ${\rm supp}(I)\subseteq{\rm supp}(\tilde{I})$, and the equality between these supports holds. ∎
Proposition 4.7 extends to the following corollary, whose proof is immediate from the definition of exact matchings (see Definition 3.1 above).
Corollary 4.10.
Let ${\mathbb{M}}$ be an interval decomposable multi-parameter persistence module, whose interval summands all satisfy the assumptions of Proposition 4.7. Let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}:=\mathtt{MMA}({\mathbb{M}},L_{\delta}(K^{2\delta}),\sigma)$ be the multi-parameter persistence module computed by $\mathtt{MMA}{}$. Then, one has
@@
4.4 Stability properties of MMA
In the previous sections, we have seen that MMA provides candidate decompositions that approximate well, and sometimes exactly, the interval decomposition of the true underlying module, when it exists. In this section, we focus on the general case, when ${\mathbb{M}}$ does not decompose into simple summands, so that there is no clear notion of induced matching $\sigma_{\mathbb{M}}$. In this case, using matching functions that pair bars that are compatible still preserves structural properties of ${\mathbb{M}}$ with MMA. Thus we start by defining compatibility.
Definition 4.11 (Compatible bars).
Let ${\mathbb{M}}$ be a multi-parameter persistence module, and let $l_{1},l_{2}\subseteq\mathbb{R}^{n}$ be two diagonal lines that are at distance $\delta$ from each other. Assume ${\rm supp}({\mathbb{M}})\cap l_{1}$ and ${\rm supp}({\mathbb{M}})\cap l_{2}$ are not empty, and let $[b^{\mathbb{M}}_{l_{1}},d^{\mathbb{M}}_{l_{1}}]$ and $[b^{\mathbb{M}}_{l_{2}},d^{\mathbb{M}}_{l_{2}}]$ be bars in ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l_{1}}})$ and ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l_{2}}})$, characterized by their endpoints. These bars are compatible if the rectangles $R_{b^{\mathbb{M}}_{l_{1}},b^{\mathbb{M}}_{l_{2}}}$, $R_{b^{\mathbb{M}}_{l_{2}},b^{\mathbb{M}}_{l_{1}}}$, $R_{d^{\mathbb{M}}_{l_{1}},d^{\mathbb{M}}_{l_{2}}}$ and $R_{d^{\mathbb{M}}_{l_{2}},d^{\mathbb{M}}_{l_{1}}}$ are flat. Moreover, we say that $[b^{\mathbb{M}}_{l_{1}},d^{\mathbb{M}}_{l_{1}}]$ is compatible with the empty set in $l_{2}$ if $\left\lVert b^{\mathbb{M}}_{l_{1}}-d^{\mathbb{M}}_{l_{1}}\right\rVert_{\infty}\leq 2\delta$.
Note that the vineyards algorithm provides compatible matching functions when $\delta$ is small enough (see Section 5.2). The main motivation for using compatible matching functions is that the candidate decompositions they induce with MMA preserve the barcodes on any diagonal line intersecting $K$ up to $2\delta$.
Proposition 4.12.
Let ${\mathbb{M}}$ be a multi-parameter persistence module. Let $\sigma$ be a compatible matching function for ${\mathcal{F}}{\mathcal{B}}({\mathbb{M}})_{L}$, and let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}:=\mathtt{MMA}({\mathbb{M}},L_{\delta}(K^{\delta}),\sigma)$. Let $l$ be a diagonal line intersecting $K$. Then, one has ${d_{\rm B}}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}),{\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}))=0$ if $l\in L$, and ${d_{\rm B}}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}),{\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}))\leq 2\delta$ otherwise.
@@ 4.4 Stability properties of MMA
Since the proof of Proposition 4.3 above is based on repeated applications of Lemma 3.9, the exact same arguments can be used to prove the first part of Proposition 4.12 (the fact that MMA with compatible matchings provides $\delta$-candidate decompositions, i.e., preserve exactly the $L$-fibered barcodes). Indeed, compatible matchings also guarantee (by definition) the flatness of rectangles created by consecutive endpoints. The general upper bound for any diagonal line is then a simple consequence of Lemma 3.10. Indeed, given a line $l\not\in L$, there must be a line $l^{*}\in L$ such that $l^{*}=l+\overrightarrow{u}$ with $\left\lVert\overrightarrow{u}\right\rVert\leq\delta$ since $L$ fills $K$. Then, one has ${d_{\rm B}}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}),{\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}))\leq{d_{\rm B}}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}),{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{*}}}))+{d_{\rm B}}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{*}}}),{\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{*}}}))+{d_{\rm B}}({\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{*}}}),{\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}))\leq\delta+0+\delta=2\delta$ by Lemma 3.10. ∎
While this result is weaker than preserving the rank invariant (which would require controlling the barcodes of all lines with positive slopes), preserving the barcodes of all diagonal lines is still a nice guarantee that ensures that the candidate decompositions provided by MMA are still appropriate, stable invariants in the general case when computed with compatible matchings. Note also that if ${\mathbb{M}}$ is interval decomposable, then compatible matchings commute with the induced matching $\sigma_{\mathbb{M}}$ under specific conditions (see Section 5.1), and can thus also be used for applying Proposition 3.4.
However, it is not possible to control more powerful distances such as ${d_{\rm I}}$ (see Section 4.5), which is expected when building visual, interpretable invariants from modules that do not decompose simply. Hence, we end this section by ensuring that MMA can still remain stable w.r.t. to the data itself by appropriately choosing the matching functions. Indeed, given two multi-parameter filtrations computed from the sublevel sets of functions $f,g$, Equation (1) ensures that the bottleneck distances between barcodes in the fibered barcodes of ${\mathbb{M}}(f)$ and ${\mathbb{M}}(g)$ are upper bounded by $\left\lVert f-g\right\rVert_{\infty}$. This in turns means that we can fix an (arbitrary) matching function $\sigma_{f}$ for computing a candidate decomposition of $f$ with MMA, and define another one $\sigma_{g}$ that commutes with $\sigma_{f}$ and the optimal partial matching $\nu$ given by those bottleneck distances. Doing so leads to the following proposition.
Proposition 4.13.
Let $f,g:S\to\mathbb{R}^{n}$ be two continuous functions defined on a simplicial complex $S$. Let $K$ be a hyperrectangle containing the minimal Betti numbers of ${\mathbb{M}}(f)$ and ${\mathbb{M}}(g)$. Let $\sigma_{f}$ be an arbitrary matching function. Then, there exists a matching function $\sigma_{g}$, and corresponding candidates $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f):=\mathtt{MMA}({\mathbb{M}}(f),L_{\delta}(K^{\delta}),\sigma_{f})$ and $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g):=\mathtt{MMA}({\mathbb{M}}(g),L_{\delta}(K^{\delta}),\sigma_{g})$ such that:
${d_{\rm I}}(\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f),\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g))\leq{d_{\rm B}}(\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f),\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g))\leq\left\lVert f-g\right\rVert_{\infty}+\delta.$
Proof.
[…]
where $\nu_{l,{d_{\rm B}}}$ denotes the optimal partial matching induced by ${d_{\rm B}}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}(f)\vphantom{\big{|}}\right|_{l}}),{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}(g)\vphantom{\big{|}}\right|_{l}}))$.
Let $I_{f},I_{g}$ denote two interval summands of $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(f)$ and $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}(g)$ that are paired by $\nu_{\cdot,{d_{\rm B}}}$. Then, controlling the bottleneck distance between these outputs of MMA simply amounts to controlling the Hausdorff distance between ${\rm supp}(I_{f})$ and ${\rm supp}(I_{g})$. In order to control this distance, let $B_{f}$ and $B_{g}$ denote the bars that induced these interval summands as per Algorithm 2. Then, one has $B_{f}\subseteq{\rm supp}(I_{f})$ and $B_{g}\subseteq B_{f}^{\gamma}$, where $\gamma=\left\lVert f-g\right\rVert_{\infty}$. Thus, since ${\rm supp}(I_{g})\subseteq B_{g}^{\delta}$ (by construction), one has ${\rm supp}(I_{g})\subseteq B_{f}^{\gamma+\delta}\subseteq{\rm supp}(I_{f})^{\gamma+\delta}$. The result follows by symmetry of $f$ and $g$. ∎
4.5 Instability of MMA w.r.t. interleaving distance
While using compatible matchings helps controlling the $L$-fibered barcodes, it is unfortunately not sufficient for bounding the interleaving distance: outputs of MMA can be very far in terms of ${d_{\rm I}}$ while the modules they are computed from are not. We provide two multi-parameter persistence modules in Figure 10 that illustrate such lack of stability. In this figure, the two bi-filtrations only differ on the middle edge (in blue) of the simplicial complex. When the appearance of this edge is delayed (as is the case for the multi-parameter filtration displayed on top), the bars in the barcodes corresponding to the lower and upper cycles of the simplicial complex get paired with the compatible matching, and create together the large red summand. On the other hand, this does not happen for the other multi-parameter filtration: the bars corresponding to these cycles are never paired and form distinct summands with same size.
Figure 10: (Left) Two bi-filtrations whose corresponding multi-parameter persistence modules in homology dimension $1$ are $\varepsilon$-interleaved. (Right) The two corresponding, significantly different interval decompositions obtained with MMA computed with a compatible matching. Intervals in these decompositions are displayed with red and yellow colors.
Figure 10 illustrates the price to pay for designing interpretable decompositions (i.e., such that each summand corresponds to a cycle of the simplicial complex) when several choices are possible: there are several different ways to assign cycles to summands in the non-interval decomposable module displayed on top of Figure 10—the one computed by MMA (shown in the figure) being one possibility. An important conjecture of this article, that is left for future work, is that stacking the candidate decompositions produced by MMA for all possible matching functions induces a complete invariant.
5 Matching functions for MMA
In this section, we study general properties of matching functions that can be used as inputs for MMA. One might wonder whether the usual distances between barcodes, such as the bottleneck or Wasserstein distances, could be used to define exact matching functions. Indeed, a major advantage of, e.g., Wasserstein distances, is that their associated matching functions are usually unique. However, when the space $\delta$ between two lines is too large, Wasserstein distances can still fail to match bars exactly, as shown in Figure 11.
Figure 11: Example interval decomposable multi-parameter persistence module with two interval summands (green and purple), and its barcodes along two lines (here the two couples of red-blue bars). Any matching function induced by, e.g., Wasserstein distances between the barcodes, will match the first red bar with the second red bar and the first blue bar with the second blue bar; however, this matching is not exact.
We show in Section 5.1 that matching functions that are compatible (see Definition 4.11) commute with the matching functions induced by interval decomposable modules. Then, we discuss matching functions induced by the vineyards algorithm in Section 5.2.
5.1 Compatible matching functions
In this section, we show that compatible matching functions commute with the induced matching function $\sigma_{{\mathbb{M}}}$ (when ${\mathbb{M}}$ is interval decomposable) under some specific conditions. In the rest of this section, given a matching function $\sigma$ that commutes with $\sigma_{\mathbb{M}}$, we will also call $\sigma$ exact for the sake of simplicity. Before stating the main result, we first show a technical lemma that relates the locations of endpoints of compatible bars with each other.
Lemma 5.1.
Let $l_{1}$ and $l_{2}$ be two $\delta$-consecutive lines, and let $[b_{1},d_{1}]:={\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{1}}})$ be the bar of an interval module along $l_{1}$. Let $[b_{2},d_{2}]$ be a bar along $l_{2}$ that is compatible with $[b_{1},d_{1}]$. Then, $d_{2}$ (resp. $b_{2}$) is included in a segment of size $\delta$ in $l_{2}$.
Proof.
Applying Lemma 3.10, one has
$d_{2}\in C:=\big{[}B_{\delta}(d_{1})\cap l_{2}\big{]}\backslash\big{[}\left\{z\in\mathbb{R}^{n}:z>d_{1}\right\}\cup\left\{z\in\mathbb{R}^{n}:z<d_{1}\right\}\big{]}$
@@ 5.1 Compatible matching functions
Since $C$ is a nonempty, totally ordered set, we can define $y:=\min C$. By construction, there exists a dimension $i$ such that $y_{i}\geq(d_{1})_{i}$, and thus $C$ must be included in the segment $[y,y+\delta\cdot\mathbf{1}]$ along $l_{2}$.
@@
Since bars that are matched under an exact matching function are always compatible, one way to construct an exact matching function between two barcodes is therefore to isolate, among all possible matching functions, the ones such that matched bars are compatible. If this family contains a single element, it must be the exact matching we are looking for. This typically happens for interval decomposable multi-parameter persistence module whose summands are sufficiently separared, as we show in the proposition below.
Proposition 5.2.
Let ${\mathbb{M}}=\bigoplus_{I\in\mathcal{I}}I$ be an interval decomposable multi-parameter persistence module. Let $\delta>0$, and $I,I^{\prime}$ be two interval summands in the decomposition of ${\mathbb{M}}$. Assume that the two following properties are satisfied:
@@
    Let $l\subset\mathbb{R}^{n}$ be a diagonal line such that ${\rm supp}(I)\cap l\neq\varnothing$ and ${\rm supp}(I^{\prime})\cap l\neq\varnothing$. Then, one has either $\left\lVert b_{l}^{I}-b_{l}^{I^{\prime}}\right\rVert_{\infty}>\delta$ or $\left\lVert d_{l}^{I}-d_{l}^{I^{\prime}}\right\rVert_{\infty}>\delta$. In other words, the endpoints of the bar in ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l}})$ and of the bar in ${\mathcal{B}}({\left.\kern-1.2ptI^{\prime}\vphantom{\big{|}}\right|_{l}})$ are at distance at least $\delta$.
@@
    The bars of length at most $2\delta$ in $I$ and $I^{\prime}$ are at distance at least $\delta$, i.e., if we let
    $S^{I}:=\left\{l:l\cap{\rm supp}(I)\neq\varnothing,\left\lVert b_{l}^{I}-d_{l}^{I}\right\rVert_{\infty}\leq 2\delta\right\},$
    (and similarly for $I^{\prime}$), one has $d_{\infty}(S^{I},S^{I^{\prime}})>\delta/2.$ In other words, a small bar in $I$ cannot be too close to a small bar in $I^{\prime}$.
Then, the matching function $\sigma_{\rm comp}$, induced by matching bars that are compatible together, is well-defined and exact.
See Figure 12 for an illustration of assumptions (1) and (2).
@@
Let $I$ and $I^{\prime}$ be two interval summands in the decomposition of ${\mathbb{M}}$. Let $l_{1}$ and $l_{2}$ be two $\delta$-consecutive lines of $L$, and let $b:={\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{1}}})$ be the bar corresponding to $I$ along $l_{1}$. We will show that $\sigma_{\rm comp}$ must match $b$ to either $b^{\prime}:={\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{2}}})$ if ${\rm supp}(I)\cap l_{2}\neq\varnothing$, or the empty set if ${\rm supp}(I)\cap l_{2}=\varnothing$.
@@
    If ${\rm supp}(I)\cap l_{2}=\varnothing$, then by Lemma 3.10, the length of $b$ is at most $\delta$, i.e., $\left\lVert b^{I}_{l_{1}}-d^{I}_{l_{1}}\right\rVert_{\infty}\leq\delta$. It is thus compatible with the empty set. Now, since $d_{\infty}(l_{1},l_{2})=\delta/2$ and since $l_{1}\in S^{I}$, assumption (2) ensures that the bar $b^{\prime\prime}:={\mathcal{B}}({\left.\kern-1.2ptI^{\prime}\vphantom{\big{|}}\right|_{l_{2}}})$ (if it exists) must be of length at least $2\delta$. In particular, it is not compatible with $b$, hence $\sigma_{\rm comp}$ cannot match $b$ to $b^{\prime\prime}$, and must match $b$ to the empty set.
@@
    If ${\rm supp}(I)\cap l_{2}\neq\varnothing$, then the bar $b^{\prime}=[b^{I}_{l_{2}},d^{I}_{l_{2}}]$ in ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{2}}})$ is compatible with $b$, as per Lemma 3.9. According to Lemma 5.1, it follows that the birthpoint and deathpoint of any bar along $l_{2}$ that is compatible to $l_{1}$ must belong to segments $s_{b},s_{d}$ of length $\delta$ that contain $b^{I}_{l_{2}}$ and $d^{I}_{l_{2}}$ respectively. Let $b^{\prime\prime}:=[b^{I^{\prime}}_{l_{2}},d^{I^{\prime}}_{l_{2}}]$ be the bar in ${\mathcal{B}}({\left.\kern-1.2ptI^{\prime}\vphantom{\big{|}}\right|_{l_{2}}})$ (if it exists). According to assumption (1), we either have $\left\lVert b_{l_{2}}^{I}-b_{l_{2}}^{I^{\prime}}\right\rVert_{\infty}>\delta$ or $\left\lVert d_{l_{2}}^{I^{\prime}}-d_{l_{2}}^{I^{\prime}}\right\rVert_{\infty}>\delta$. In particular this means that either $b^{I^{\prime}}_{l_{2}}\not\in s_{b}$ or $d^{I^{\prime}}_{l_{2}}\not\in s_{d}$. Hence $b^{\prime\prime}$ is not compatible with $b$, and $\sigma_{\rm comp}$ must match $b$ to $b^{\prime}$.
In both cases, $\sigma_{\rm comp}$ is well-defined and exact. ∎
5.2 Vineyards matching functions
This section is dedicated to the vineyards algorithm for simplicial complexes. It is quite technical and can be skipped by readers who are most interested in the general exposition. The vineyards algorithm provides a matching function $\sigma_{\rm vine}$ between barcodes, so the purpose of this section is to show that it is exact, i.e., that it commutes with the induced matching function $\sigma_{\mathbb{M}}$, provided that ${\mathbb{M}}$ is interval decomposable. Since vineyards are heavily based on simplicial homology, we first recall the basics of persistent homology from simplicial complexes in Section 5.2.1. Then, we provide an analysis of the vineyards algorithm in Section 5.2.2.

### v4

@@ 5.2 Exact reconstruction
    Assume $S\not\subseteq L[I]$. In that case, there must be a facet of codirection $j$, for some $j\in\mathcal{J}^{-}$, that intersects $S$. Since one has $c^{-}_{j}\leq(c^{-}+\overrightarrow{v})_{j}<\alpha_{j}$ for any $j\in\mathcal{J}^{-}$, this means that the facet would not intersect $K$, which yields to a contradiction as per Definition 24, item (4).
This concludes that ${I}\subseteq{\tilde{I}}$, and the equality between these supports holds. ∎
Lemma 9 extends to the following proposition, whose proof is immediate from the definition of induced matchings.
Proposition 12 (Exact recovery).
Let $\mathbb{M}$ be a f.p. interval decomposable $n$-parameter persistence module. Let $K$ be a rectangle in $\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}$, and $L:=L_{\delta}(K^{2\delta})$ be the $\delta$-grid of lines of the offset $K^{2\delta}$. Assume that all interval summands of $\mathbb{M}$ are $\delta$-discretely presented, and let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}:=\mathtt{MMA}{}({\mathbb{M}},L,\sigma)$}, where $\sigma$ is a matching function that commutes with the induced matching function $\sigma_{\mathbb{M}}$. Then, one has:
@@
Note that f.p. interval decomposable modules are always made of $\delta$-discretely presented interval summands, for small enough yet positive $\delta$ (one can take for instance the smallest distance $\delta_{\rm exact}>0$ between two distinct graded Betti numbers).
One might wonder whether the usual distances between barcodes, such as the bottleneck or Wasserstein distances, could be used to define matching functions that commute with induced matching functions. Indeed, a major advantage of, e.g., Wasserstein distances, is that their associated matching functions are usually unique. However, when the space $\delta$ between two lines is too large, the matching functions induced by Wasserstein distances can still fail to be induced, as shown in Figure 12. In the next section, we discuss how to design such matching functions from compatible matching functions.
Figure 12: Example of interval decomposable module with two interval summands (green and purple), and its barcodes along two lines (here the two couples of red-blue bars). Any matching function induced by, e.g., Wasserstein distances between the barcodes, will match the first red bar with the second red bar and the first blue bar with the second blue bar; however, this matching is not induced.
6 Finding compatible and induced matching functions
In this section, we discuss how to design matching functions that are compatible or that commute with induced matching functions, so as to satisfy the assumptions of Proposition 3, Proposition 5 and Proposition 12. In particular, in Section 6.1 we restrict to interval decomposable modules, and we show that compatible and usual matching functions always commute with induced matching functions for small enough $\delta$ and generic assumptions. Then, in Section 6.2, we show that the vineyards algorithm induces compatible matching functions when $n=2$.
6.1 Induced matching functions for interval decomposable modules
In this section, we show, given some interval decomposable module $\mathbb{M}$, that compatible matching functions commute with the induced matching function $\sigma_{{\mathbb{M}}}$ under some specific conditions. In the rest of this section, given a matching function $\sigma$ that commutes with $\sigma_{\mathbb{M}}$, we will also call $\sigma$ induced for the sake of simplicity. Before stating the main result, we first show a technical lemma that relates the locations of endpoints of compatible bars with each other.
Lemma 13.
Let $l_{1}$ and $l_{2}$ be two $\delta$-consecutive lines, and let $b_{1},d_{1}):={\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)$ be the bar of a f.p. interval module along $l_{1}$. Let $b_{2},d_{2})$ be a bar along $l_{2}$ that is compatible with $b_{1},d_{1})$. Then, $d_{2}$ (resp. $b_{2}$) is included in a segment of size $\delta$ in $l_{2}$.
@@ 6.1 Induced matching functions for interval decomposable modules
Applying Lemma 2, one has:
$d_{2}\in C:=\big[B_{\delta}(d_{1})\cap l_{2}\big]\backslash\big[\left\{z\in\mathbb{R}^{n}:z>d_{1}\right\}\cup\left\{z\in\mathbb{R}^{n}:z<d_{1}\right\}\big].$
@@
Since $C$ is a nonempty, totally ordered set, we can define $y:=\min C$. By construction, there exists a dimension $i$ such that $y_{i}\geq(d_{1})_{i}$, and thus $C$ must be included in the segment $y,y+\delta\cdot\mathbf{1}]$ along $l_{2}$.
@@
Since bars that are matched under an induced matching function are always compatible, one way to construct an induced matching function between two barcodes is therefore to isolate, among all possible matching functions, the ones such that matched bars are compatible. If this family contains a single element, it must be the induced matching we are looking for. This typically happens for interval decomposable multi-parameter persistence modules whose summands are sufficiently separared, as we show in the proposition below.
Proposition 14.
Let ${\mathbb{M}}=\bigoplus_{I\in\mathcal{I}}{k^{I}}$ be a f.p. interval decomposable $n$-parameter persistence module. Let $\delta>0$, and ${k^{I}},{k^{I^{\prime}}}$ be two interval summands in the decomposition of $\mathbb{M}$}. Assume that the two following properties are satisfied:
@@
    Let $l\subset\mathbb{R}^{n}$ be a diagonal line such that ${I}\cap l\neq\varnothing$ and ${I^{\prime}}\cap l\neq\varnothing$.
    Then, one has either $\left\lVert b_{l}^{I}-b_{l}^{I^{\prime}}\right\rVert_{\infty}>\delta$ or $\left\lVert d_{l}^{I}-d_{l}^{I^{\prime}}\right\rVert_{\infty}>\delta$. In other words, the endpoints of the bar in ${\mathcal{B}}\left({{\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l}}}\right)$ and of the bar in ${\mathcal{B}}\left({{\left.\kern-1.2ptk^{I^{\prime}}\vphantom{\big|}\right|_{l}}}\right)$ are at distance at least $\delta$.
@@
    The bars of length at most $2\delta$ in $I$ and $I^{\prime}$ are at distance at least $\delta$, i.e., if we let:
    $S^{I}:=\left\{l:l\cap{I}\neq\varnothing,\left\lVert b_{l}^{I}-d_{l}^{I}\right\rVert_{\infty}\leq 2\delta\right\},$
    (and similarly for $I^{\prime}$), one has
    $d_{\infty}\left(S^{I},S^{I^{\prime}}\right)>\delta/2.$
    In other words, a small bar in $I$ cannot be too close to a small bar in $I^{\prime}$.
Then, the matching function $\sigma_{\rm comp}$, induced by matching bars that are compatible together, is well-defined and induced from $\mathbb{M}$.
Note that, upon using chunk reduction (36) and infinitesimal perturbations, or whenever the graded Betti numbers of $\mathbb{M}$ are independent and identically distributed from a non-singular distribution, it is always possible to ensure that Assumptions (1) and (2) are satisfied for a given f.p. interval decomposable module $\mathbb{M}$ and small enough $\delta$ (see also the paragraph after Proposition 12). See Figure 13 for an illustration of Assumptions (1) and (2).
@@
Let $k^{I}$ and $k^{I^{\prime}}$ be two interval summands in the decomposition of $\mathbb{M}$}. Let $l_{1}$ and $l_{2}$ be two $\delta$-consecutive lines of $L$, and let $b:={\mathcal{B}}\left({\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l_{1}}}\right)$ be the bar corresponding to $I$ along $l_{1}$. We will show that $\sigma_{\rm comp}$ must match $b$ to either $b^{\prime}:={\mathcal{B}}\left({\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l_{2}}}\right)$ if ${I}\cap l_{2}\neq\varnothing$, or the empty set if ${I}\cap l_{2}=\varnothing$.
@@
    If ${\rm supp}(I)\cap l_{2}=\varnothing$, then by Lemma 2, the length of $b$ is at most $\delta$, i.e., $\left\lVert b^{I}_{l_{1}}-d^{I}_{l_{1}}\right\rVert_{\infty}\leq\delta$.
    It is thus compatible with the empty set. Now, since $d_{\infty}(l_{1},l_{2})=\delta/2$ and since $l_{1}\in S^{I}$, Assumption (2) ensures that the bar $b^{\prime\prime}:={\mathcal{B}}\left({\left.\kern-1.2ptk^{I^{\prime}}\vphantom{\big|}\right|_{l_{2}}}\right)$ (if it exists) must be of length at least $2\delta$. In particular, it is not compatible with $b$, hence $\sigma_{\rm comp}$ cannot match $b$ to $b^{\prime\prime}$, and must match $b$ to the empty set.
@@
    If ${\rm supp}(I)\cap l_{2}\neq\varnothing$, then the bar $b^{\prime}=[b^{I}_{l_{2}},d^{I}_{l_{2}})$ in ${\mathcal{B}}\left({\left.\kern-1.2ptk^{I}\vphantom{\big|}\right|_{l_{2}}}\right)$
    is compatible with $b$, as per Lemma 1. According to Lemma 13, it follows that the birthpoint and deathpoint of any bar along $l_{2}$ that is compatible to $l_{1}$ must belong to segments $s_{b},s_{d}$ of length $\delta$ that contain $b^{I}_{l_{2}}$ and $d^{I}_{l_{2}}$ respectively. Let $b^{\prime\prime}:=[b^{I^{\prime}}_{l_{2}},d^{I^{\prime}}_{l_{2}})$ be the bar in ${\mathcal{B}}\left({\left.\kern-1.2ptk^{I^{\prime}}\vphantom{\big|}\right|_{l_{2}}}\right)$ (if it exists).
    According to Assumption (1), we either have $\left\lVert b_{l_{2}}^{I}-b_{l_{2}}^{I^{\prime}}\right\rVert_{\infty}>\delta$ or $\left\lVert d_{l_{2}}^{I^{\prime}}-d_{l_{2}}^{I^{\prime}}\right\rVert_{\infty}>\delta$. In particular this means that either $b^{I^{\prime}}_{l_{2}}\not\in s_{b}$ or $d^{I^{\prime}}_{l_{2}}\not\in s_{d}$. Hence $b^{\prime\prime}$ is not compatible with $b$, and $\sigma_{\rm comp}$ must match $b$ to $b^{\prime}$.
In both cases, $\sigma_{\rm comp}$ is well-defined and induced from $\mathbb{M}$. ∎
One can check that the proof of Proposition 14 extends easily to the matching functions associated to the Wasserstein distances and the vineyards algorithm. Indeed, their associated matching functions are unique when $\delta$ becomes small enough, and thus must correspond to the only compatible matching $\sigma_{\rm comp}$ identified in Proposition 14.
6.2 The vineyards algorithm for general $2$-parameter modules
In this section, we show that the matching function associated to the vineyards algorithm for simplicial complexes is compatible, for small enough $\delta$ and $n=2$. This section is quite technical and can be skipped by readers who are most interested in the general exposition. Since vineyards are heavily based on simplicial homology, we first recall the basics of persistent homology from simplicial complexes in Section 6.2.1. Then, we provide an analysis of the vineyards algorithm in Section 6.2.2.